OpenStudy (anonymous):

Ok I have a couple more I can't get. These are not long problems though. Thank God. 1/9x^2+11/3x=80

OpenStudy (anonymous):

3x/9x^2+11=80

OpenStudy (anonymous):

3x= 80(9x^2+11)

OpenStudy (anonymous):

Lova is it 1 11 ----- x^2 + --- x = 80 9 3

OpenStudy (anonymous):

The problem reads: Estimate a safe speed x for a car for a stopping distance of 80 feet. Use the formula below

OpenStudy (anonymous):

If my equation is correct, then multiply every term by 9 because that is the common factor

OpenStudy (anonymous):

you should get x^2 + 33x = 270

OpenStudy (anonymous):

I put that in my assignment and it says that is not right. I have been trying to get this on for a while now

myininaya (myininaya):

9(80)=720 right?

OpenStudy (dumbcow):

blexting i think you had a typo* x^2 + 33x = 720 move 720 over and try to factor

OpenStudy (anonymous):

OH MY you are right... x^2 + 33x = 720

OpenStudy (anonymous):

x^2 + 33x - 720 = 0

myininaya (myininaya):

720=10(72)=20(36)=60(12)=30(24)=15(48) -720=48(-15) 33=48+(-15)

OpenStudy (anonymous):

(x + 48)(x - 15) = 0 x = -48 or x = 15 Since this deals with speed use the x = 15

OpenStudy (anonymous):

thanks myininaya

myininaya (myininaya):

your welcome

OpenStudy (anonymous):

You guys are awesome. Thank you so much

OpenStudy (amistre64):

yay!!